# Survey on Publicly Available Sinhala Natural Language Processing Tools and Research

### v23

@@ 3.3 Dictionaries
A necessary component for the purpose of bridging Sinhala and English resources are English-Sinhala dictionaries. The earliest and most extensive Sinhala-English dictionary available for consumption was by Malalasekera [155]. However, this dictionary is locked behind copyright laws and is not available for public research and development. This copyright issue is shared with other printed dictionaries [156, 157, 158, 159, 160, 161] as well. The dictionary by Kulatunga [162] is publicly available for usage through an online web interface but does not provide API access or means to directly access the data set. The largest publicly available English-Sinhala dictionary data set is from a discontinued FireFox plug-in EnSiTip [163] which bears a more than passing resemblance to the above dictionary by Kulatunga [162]Hettige and Karunananda [74] claim to to have created a lexicon to help in their attempt to create a system capable of English-to-Sinhala machine translation. A review on the requirements for English-Sinhala smart bilingual dictionary was conducted by Samarawickrama and Hettige [164]. The study by Wickramasinghe and De Silva [140] introduced a large Sinhala-English dictionary which has $1,368,416$ unfiltered and $195,255$ filtered En-Sipairs. Both their code and the two versions of the dictionary are publicly available on github [^91] https://github.com/kasunw22/sinhala-para-dict/tree/main. A Sinhala-Sinhala dictionary named Wahara [^92] http://crawler.nlpc.uom.lk/ was built and is publicly maintained by the University of Moratuwa.
@@ 3.4 WordNets
WordNets [167] are extremely powerful and act as a versatile component of many NLP applications. They encompass a number of linguistic properties which exist between the words in the lexicon of the language including but not limited to: hyponymy, hypernymy, synonymy, and meronymy. Their uses range from simple gazetteer listing applications [32] to information extraction based on semantic similarity [168, 169] or semantic oppositeness [170]. An attempt has been made to build a Sinhala Wordnet by Wijesiri et al. [43]. For a time it was hosted on [171] but it too is now defunct and all the data and applications are lost other than what Arukgoda et al. [172] have cloned to use in their application uploaded on github [^93] https://github.com/jseanm1/aruthSWSD. However, even at its peak, due to the lack of volunteers for the crowd-sourced methodology of populating the WordNet, it was at best an incomplete product. Another effort to build a Sinhala Wordnet was initiated by Welgama et al. [173] independently from above; but it too have stopped progression even before achieving the completion level of Wijesiri et al. [43].
@@ 3.5 Morphological Analyzers
The next attempt by Herath et al. [174] creates a modular unit structure for morphological analysis of Sinhala. Much later, as a step on their efforts to create a system with the ability to do English-to-Sinhala machine translation, Hettige and Karunananda [71] claim to have created a morphological analyzer (void of any public data or code), which links to their studies of a Sinhala parser [72] and computational grammar [20]Hettige et al. [83] further propose a multi-agent System for morphological analysis. Welgama et al. [175] attempted to evaluate machine learning approaches for Sinhala morphological analysis. Yet another independent attempt to create a morphological parser for Sinhala verbs was carried out by Fernando and Weerasinghe [176]. Later, another study, which was restricted to morphological analysis of Sinhala verbs was conducted by Dilshani and Dias [177]. There was no indication on whether this work was continued to cover other types of words. Further, other than this singular publication, no data or tools were made publicly accessible. Nandathilaka et al. [178] proposed a rule based approach for Sinhala lemmatizing. The work by Welgama et al. [179] claim to have set a set of gold standard definitions for the morphology of Sinhala Words; but given that their results are not publicly available, further usage or confirmation of these claims cannot not be done. The table V provides a comparative summery of the discussion above. The combined study introduced a rule-based stemmer [180] and a tokenizer [181] for Sinhala. A later work by Kumarasinghe et al. [182] named SinMorphy used a comprehensive vocabulary of Sinhala words to conduct rule-based morphological analysis on Sinhala. Ekanayaka et al. [183] compared the effectiveness of using RNN, LSTM, and GRU for Sinhala morphological analysis of Sinhala and Sinhlish deatasets and found the BiGRU gives the highest accuracy of 87.96%.
@@ 3.6 Part of Speech Taggers
Within a single group which did one of the above stochastic studies [185], yet another set of studies was carried out to create a Sinhala PoS tagger starting with the foundation of Jayaweera and Dias [187] which then extended to a Hidden Markov Model (HMM) based approach [188] and an analysis of unknown words [189, 190]. Further, this group presented a comparison of few Sinhala PoS taggers that are available to them [191]. A RESTFul PoS tagging web service created by Jayaweera and Dias [192] using the above research can still be accessed [^94] http://bit.ly/2F0jKid via POST and GET. A hybrid PoS tagger for Sinhala language was proposed by Gunasekara et al. [193]. The study by Kothalawala et al. [194] discussed the data availability problem in NLP with a Sinhala POS tagging experiment among others.  Withanage and Silva [195] proposed a stochastic POS tagger based on a small 10,000 word corpus drawn from Facebook and Twitter.  Wijerathna [196] used Support Vector Machines (SVM) to tag Sinhala text with 30 the PoS tags that were proposed by Fernando et al. [111]. The study by Sathsarani et al. [197] compared rule-based and stochastic models against deep learning models in the task of Sinhala PoS tagging.
@@ 3.8 Named Entity Recognition Tools
The table VI provides a comparative summary of the discussion above. It should be noted that all three of these models only tag NEs of types: person names, location names and organization names. The Ananya system by Manamini et al. [114] is available to download at GitHub  [^95] http://bit.ly/2XrwCoK. The data and code for the approaches by Dahanayaka and Weerasinghe [202] and by Senevirathne et al. [203] are not accessible to the public. Azeez and Ranathunga [204] proposed a fine-grained NER model for Sinhala building on their earlier work on NER [114] and PoS tagging [113]Anuruddha [205] proposed a method based on reinforcement learning for Sinhala NER. A Sinhala NER system restricted to the sports domain was introduced by Wijesinghe and Tissera [206], where they attempted to utilise classical machine learning models. The work by Mallikarachchi et al. [207] used support vector machines to detect Sinhala named entities of types person, location and, organization. Even though the main task reported in the study by [208] is clustering Sinhala news articles, they used NER tools to do so. The work by Gurgurov et al. [209] compared the performance of LLMs for the Sinhala NER task. The highest results are obtained for this task in the Masked Language Modeling objective for Sinhala. They have used a mixture of ConceptNet (CN) and/or Wikipedia (wiki) language adapters in conjunction with Task Adapters (TA).
@@ 3.9 Semantic Similarity
A Sinhala semantic similarity measure has been developed for short sentences by Kadupitiya et al. [210]. This work has been then extended by Kadupitiya et al. [211] for the application use case of short answer grading. Data and tools for these projects are not publicly available. A sentence similarity measurement which uses Siamese neural networks was developed by Nilaxan and Ranathunga [212] where they demonstrate their results for Sinhala and Tamil. A cross-lingual document similarity measurement using the use-case of Sinhala and English was developed by Isuranga et al. [213].
@@ 3.10 Text Classification
Text classification is a popular application on the semantic layer of the NLP stack. A very basic Sinhala text classification using Naïve Bayes Classifier, Zipf’s Law Behavior, and SVMs was attempted by Gallege [214]. A smaller implementation of Sinhala news classification has been attempted by de Silva [24]. As mentioned in Section 3.2, their news corpus is publicly available [^96] https://osf.io/tdb84/. Another attempt at Sinhala text classification using six popular rule-based algorithms was done by Lakmali and Haddela [215]. Even though they talk about building a corpus named SinNG5, they do not indicate of means for others to obtain the said corpus. Another study by Kumari and Haddela [216] utilizes the SinNG5 corpus as the data set for their attempt to use LIME [217] for human interpretability of Sinhala document classification. However, they too do not provide access to the corpus. Nanayakkara and Ranathunga [218] have implemented a system which uses corpus-based similarity measures for Sinhala text classification.  Gunasekara and Haddela [219, 220] claim to have created a context-aware stop word extraction method for Sinhala text classification based on simple TF-IDF. An LSTM based textual entailment system for Sinhala was proposed by Jayasinghe and Sirts [221]Demotte and Ranathunga [222] proposed a dual-state capsule network architecture for text classification where they demonstrated their methodology on the Sinhala data set established by Senevirathne et al. [223]. The attempt by Sameemdeen and Selvanthan [224] considers three classical machine learning algorithms (Naïve Bayes,SVM,and KNN) and then goes on to briefly discusses the pros and cons of previous attempts by: Nanayakkara and Ranathunga [218]Gunasekara and Haddela [219]Lakmali and Haddela [215], and Buddhika et al. [225]. Then they propose active learning [226, 227] as an alternative. However, no experimental results of how active learning would improve Sinhala text classification are given. The study by Bandara et al. [228] proposed and ontology-based approach for Sinhala fake news detection. However, their literature survey did not cover seminal papers in OBIE such as the work by Wimalasuriya and Dou [32]. This has impacted the overall methodology that was presented. A novel, domain-idependant, and domain-adaptive text classification framework named AdaptText for Sinhala text classification was proposed by Kodithuwakku and Hettiarachchi [229]. A simple TF-IDF based Sinhala text classification system was proposed by Koralage [230]. The study by [231] used Genetic Algorithm on Lucene [^97] https://lucene.apache.org/ search queries to obtain interpretable classification models for Sinhala text documents. The approach proposed by Rathnayake et al. [232] uses adapter-based [233, 234, 235, 236, 237, 238] fine-tuning on XLM-R [239], for code-mixed and code-switched text. Kirindage and Godewithana [240] used LDA [241] to develop Sinhala news topic hierarchies and categorize Sinhala news documents using the said topic hierarchies. Hettigoda [242] collected English-Sinhala code-mixed comments from Facebook pages belonging to Clothing industrial online businesses in Sri Lanka and classified them in to the classes: Inquiries, Maybe Inquiries, and Not Inquiries. A comparative analysis on BERT based models for Sinhala text classification was conducted by Dhananjaya et al. [104]. Both their code [^98] https://github.com/nlpcuom/Sinhala-text-classification and data (sin-cc-15M corpus9999993https://tinyurl.com/42un7a9y, Sinhala News source classification data set100100100https://bit.ly/3PBTyfo, Sinhala News category classification data set101101101https://bit.ly/3AsfK7u, and Sinhala writing style classification data set102102102https://bit.ly/3PARuUS) are publicly available. The study by Wijayarathna and Jayalal [243] used classical machine learning techniques including Random Forest to classify fake news in Sinhala on Twitter. Chathuranga and Ranathunga [244] used capsule-based methods recommended by Senevirathne et al. [223] to classify Sinhala-English code-mixed data. The work by Weerasiri et al. [245] compared Word2Vec, FastText and Doc2vec [246] against TF-IDF for Sinhala news document classification. The study by Caldera et al. [247] used stacked LSTM to classify Sinhala and Singlish text discussing COVID-19 on Youtube and Twitter.
@@ 3.11 Sentiment Analysis
A simple MLP-based method to classify sentiments in Sinhala text was initially proposed by Medagoda [248] based on their prior work [249]. A word2vec based tool103103103http://bit.ly/2QKI9Np for sentiment analysis of Sinhala news comments is available. A methodology for constructing a sentiment lexicon for Sinhala Language in a semi-automated manner based on a given corpus was proposed by Chathuranga et al. [250]Demotte et al. [251] proposed a sentiment analysis system based on sentence-state LSTM Networks for Sinhala news comments. In the subsequent work [252], they used word similarity to generate a Sinhala semantic lexicon. They followed this up with a further study [223] which discussed a number of other deep learning models such as RNN and Bi-LSTM in the domain of Sinhala sentiment analysis. Jayasuriya et al. [253] proposed a method to classify Sinhala posts in the domain of sports into positive and negative class sentiments. Ranathunga and Liyanage [254] claimed that using word embedding models as semantic features can compensate for the lack of well developed language-specific linguistic or language resources in the case of analysing sentiment of Sinhala news comments. Jayasuriya et al. [255] conducted a comparative study between word N-grams and character N-grams in the task of semantic classification of Sinhala content in social media. Which they soon followed up with an ensemble approach [256]. The work by Karunarathne [257] used word embedding to analyse the sentiment of manually annotated Sinhala Tweets. Abeyratne and Jayaratne [258] conducted a multi model analysis on classifying Sinhala songs by emotion.
@@
The work by Jayawickrama et al. [259, 260] used the data set released by Wijeratne and de Silva [96] to predict the reactions induced by Sinhala Facebook posts. They then extended the work [261] and compared the results obtained with their data set against that of Senevirathne et al. [223]Aththanayaka and Naleer [262] used Random forest, Support vector machines, and Multinomial Naïve Bayes models to analyse sentiment in Sinhala-English code-mixed text from social media. The ACTSEA dataset104104104https://bit.ly/3oZoziI for Sinhala sentiment analysis was introduced by Jenarthanan et al. [131]. The dataset contains 318,308 Sinhala tweets annotated with emotions. The study by Dhananjaya et al. [263] uses a sentiment lexicon of a high‐resource language to fine-tune Pre‐trained multilingual language models (PMLMs) such as mBERT and XLM‐R on an intermediate task which in turn is then used on the sentiment classification task of Sinhala. A sentiment-annotated dataset of Sinhala-English code mixed YouTube comments has been created by Uthpala and Thirukumaran [264]. However, given that even their paper is behind a paywall, there is no free and public way to access this dataset. The work by Gurgurov et al. [209] compared the performance of LLMs for the Sinhala sentiment analysis task. They report the highest results for the Masked Language Modeling objective for Sinhala in both of their configurations where they used a mixture of ConceptNet (CN) and/or Wikipedia (wiki) language adapters in conjunction with Task Adapters (TA).

### v24

@@ 3.3 Dictionaries
A necessary component for the purpose of bridging Sinhala and English resources are English-Sinhala dictionaries. The earliest and most extensive Sinhala-English dictionary available for consumption was by Malalasekera [174]. However, this dictionary is locked behind copyright laws and is not available for public research and development. This copyright issue is shared with other printed dictionaries [175, 176, 177, 178, 179, 180] as well. The dictionary by Kulatunga [181] is publicly available for usage through an online web interface but does not provide API access or means to directly access the data set. The largest publicly available English-Sinhala dictionary data set is from a discontinued FireFox plug-in EnSiTip [182] which bears a more than passing resemblance to the above dictionary by Kulatunga [181]Hettige and Karunananda [79] claim to to have created a lexicon to help in their attempt to create a system capable of English-to-Sinhala machine translation. A review on the requirements for English-Sinhala smart bilingual dictionary was conducted by Samarawickrama and Hettige [183]. The study by Wickramasinghe and De Silva [147] introduced a large Sinhala-English dictionary which has $1,368,416$ unfiltered and $195,255$ filtered En-Sipairs. Both their code and the two versions of the dictionary are publicly available on github102102102https://github.com/kasunw22/sinhala-para-dict/tree/main. A Sinhala-Sinhala dictionary named Wahara103103103http://crawler.nlpc.uom.lk/ was built and is publicly maintained by the University of Moratuwa.
@@ 3.4 WordNets
WordNets [186] are extremely powerful and act as a versatile component of many NLP applications. They encompass a number of linguistic properties which exist between the words in the lexicon of the language including but not limited to: hyponymy, hypernymy, synonymy, and meronymy. Their uses range from simple gazetteer listing applications [37] to information extraction based on semantic similarity [187, 188] or semantic oppositeness [189]. An attempt has been made to build a Sinhala Wordnet by Wijesiri et al. [48]. For a time it was hosted on [190] but it too is now defunct and all the data and applications are lost other than what Arukgoda et al. [191] have cloned to use in their application uploaded on github104104104https://github.com/jseanm1/aruthSWSD. However, even at its peak, due to the lack of volunteers for the crowd-sourced methodology of populating the WordNet, it was at best an incomplete product. Another effort to build a Sinhala Wordnet was initiated by Welgama et al. [192] independently from above; but it too have stopped progression even before achieving the completion level of Wijesiri et al. [48].
@@ 3.5 Morphological Analyzers
The next attempt by Herath et al. [193] creates a modular unit structure for morphological analysis of Sinhala. Much later, as a step on their efforts to create a system with the ability to do English-to-Sinhala machine translation, Hettige and Karunananda [76] claim to have created a morphological analyzer (void of any public data or code), which links to their studies of a Sinhala parser [77] and computational grammar [21]Hettige et al. [88] further propose a multi-agent System for morphological analysis. Welgama et al. [194] attempted to evaluate machine learning approaches for Sinhala morphological analysis. Yet another independent attempt to create a morphological parser for Sinhala verbs was carried out by Fernando and Weerasinghe [195]. Later, another study, which was restricted to morphological analysis of Sinhala verbs was conducted by Dilshani and Dias [196]. There was no indication on whether this work was continued to cover other types of words. Further, other than this singular publication, no data or tools were made publicly accessible. Nandathilaka et al. [197] proposed a rule based approach for Sinhala lemmatizing. The work by Welgama et al. [198] claim to have set a set of gold standard definitions for the morphology of Sinhala Words; but given that their results are not publicly available, further usage or confirmation of these claims cannot not be done. The table V provides a comparative summery of the discussion above. The combined study introduced a rule-based stemmer [199] and a tokenizer [200] for Sinhala. A later work by Kumarasinghe et al. [201] named SinMorphy used a comprehensive vocabulary of Sinhala words to conduct rule-based morphological analysis on Sinhala. Ekanayaka et al. [202] compared the effectiveness of using RNN, LSTM, and GRU for Sinhala morphological analysis of Sinhala and Sinhlish deatasets and found the BiGRU gives the highest accuracy of 87.96%. The comparative study by [203] compared Sinhala morphology to that of a number of other languages in hopes of making subsequent machine translation work easier.
@@ 3.6 Part of Speech Taggers
Within a single group which did one of the above stochastic studies [205], yet another set of studies was carried out to create a Sinhala PoS tagger starting with the foundation of Jayaweera and Dias [207] which then extended to a Hidden Markov Model (HMM) based approach [208] and an analysis of unknown words [209, 210]. Further, this group presented a comparison of few Sinhala PoS taggers that are available to them [211]. A RESTFul PoS tagging web service created by Jayaweera and Dias [212] using the above research can still be accessed105105105http://bit.ly/2F0jKid via POST and GET. A hybrid PoS tagger for Sinhala language was proposed by Gunasekara et al. [213]. The study by Kothalawala et al. [214] discussed the data availability problem in NLP with a Sinhala POS tagging experiment among others.  Withanage and Silva [215] proposed a stochastic POS tagger based on a small 10,000 word corpus drawn from Facebook and Twitter.  Wijerathna [216] used Support Vector Machines (SVM) to tag Sinhala text with 30 the PoS tags that were proposed by Fernando et al. [115]. The study by Sathsarani et al. [217] compared rule-based and stochastic models against deep learning models in the task of Sinhala PoS tagging.
@@ 3.8 Named Entity Recognition Tools
The table VI provides a comparative summary of the discussion above. It should be noted that all three of these models only tag NEs of types: person names, location names and organization names. The Ananya system by Manamini et al. [118] is available to download at GitHub 106106106http://bit.ly/2XrwCoK. The data and code for the approaches by Dahanayaka and Weerasinghe [222] and by Senevirathne et al. [223] are not accessible to the public. Azeez and Ranathunga [224] proposed a fine-grained NER model for Sinhala building on their earlier work on NER [118] and PoS tagging [117]Anuruddha [225] proposed a method based on reinforcement learning for Sinhala NER. A Sinhala NER system restricted to the sports domain was introduced by Wijesinghe and Tissera [226], where they attempted to utilise classical machine learning models. The work by Mallikarachchi et al. [227] used support vector machines to detect Sinhala named entities of types person, location and, organization. Even though the main task reported in the study by [228] is clustering Sinhala news articles, they used NER tools to do so. The work by Gurgurov et al. [229] compared the performance of LLMs for the Sinhala NER task. The highest results are obtained for this task in the Masked Language Modeling objective for Sinhala. They have used a mixture of ConceptNet (CN) and/or Wikipedia (wiki) language adapters in conjunction with Task Adapters (TA). A data set for Sinhala NER named multiNER107107107https://github.com/suralk/multiNER was introduced by Ranathunga et al. [168]. It includes parallel NER data for Sinhala, Tamil, and English.
@@ 3.9 Semantic Similarity
A Sinhala semantic similarity measure has been developed for short sentences by Kadupitiya et al. [230]. This work has been then extended by Kadupitiya et al. [231] for the application use case of short answer grading. Data and tools for these projects are not publicly available. A sentence similarity measurement which uses Siamese neural networks was developed by Nilaxan and Ranathunga [232] where they demonstrate their results for Sinhala and Tamil. A cross-lingual document similarity measurement using the use-case of Sinhala and English was developed by Isuranga et al. [233]. Deepal et al. [234] proposed to use Siamese network which is a hybrid of LSTM and CNN to calculate Sinhala sentence similarity. They report an F1 of 0.9041 for their model calculated by cosine distance. They discuss how they used a number of publicly available data sets and refined them to be used in their task, nevertheless, they do not publicly release the resulting data set.
@@ 3.10 Text Classification
Text classification is a popular application on the semantic layer of the NLP stack. A very basic Sinhala text classification using Naïve Bayes Classifier, Zipf’s Law Behavior, and SVMs was attempted by Gallege [235]. A smaller implementation of Sinhala news classification has been attempted by de Silva [28]. As mentioned in Section 3.2, their news corpus is publicly available108108108https://osf.io/tdb84/. Another attempt at Sinhala text classification using six popular rule-based algorithms was done by Lakmali and Haddela [236]. Even though they talk about building a corpus named SinNG5, they do not indicate of means for others to obtain the said corpus. Another study by Kumari and Haddela [237] utilizes the SinNG5 corpus as the data set for their attempt to use LIME [238] for human interpretability of Sinhala document classification. However, they too do not provide access to the corpus. Nanayakkara and Ranathunga [239] have implemented a system which uses corpus-based similarity measures for Sinhala text classification.  Gunasekara and Haddela [240, 241] claim to have created a context-aware stop word extraction method for Sinhala text classification based on simple TF-IDF. An LSTM based textual entailment system for Sinhala was proposed by Jayasinghe and Sirts [242]Demotte and Ranathunga [243] proposed a dual-state capsule network architecture for text classification where they demonstrated their methodology on the Sinhala data set established by Senevirathne et al. [244]. The attempt by Sameemdeen and Selvanthan [245] considers three classical machine learning algorithms (Naïve Bayes,SVM,and KNN) and then goes on to briefly discusses the pros and cons of previous attempts by: Nanayakkara and Ranathunga [239]Gunasekara and Haddela [240]Lakmali and Haddela [236], and Buddhika et al. [246]. Then they propose active learning [247, 248] as an alternative. However, no experimental results of how active learning would improve Sinhala text classification are given. The study by Bandara et al. [249] proposed and ontology-based approach for Sinhala fake news detection. However, their literature survey did not cover seminal papers in OBIE such as the work by Wimalasuriya and Dou [37]. This has impacted the overall methodology that was presented. A novel, domain-idependant, and domain-adaptive text classification framework named AdaptText for Sinhala text classification was proposed by Kodithuwakku and Hettiarachchi [250]. A simple TF-IDF based Sinhala text classification system was proposed by Koralage [251]. The study by [252] used Genetic Algorithm on Lucene109109109https://lucene.apache.org/ search queries to obtain interpretable classification models for Sinhala text documents. The approach proposed by Rathnayake et al. [253] uses adapter-based [254, 255, 256, 257, 258, 259] fine-tuning on XLM-R [260], for code-mixed and code-switched text. Kirindage and Godewithana [261] used LDA [262] to develop Sinhala news topic hierarchies and categorize Sinhala news documents using the said topic hierarchies. Hettigoda [263] collected English-Sinhala code-mixed comments from Facebook pages belonging to Clothing industrial online businesses in Sri Lanka and classified them into the classes: Inquiries, Maybe Inquiries, and Not Inquiries. A comparative analysis on BERT based models for Sinhala text classification was conducted by Dhananjaya et al. [107]. Both their code110110110https://github.com/nlpcuom/Sinhala-text-classification and data (sin-cc-15M corpus1111111113https://tinyurl.com/42un7a9y, Sinhala News source classification data set112112112https://bit.ly/3PBTyfo, Sinhala News category classification data set113113113https://bit.ly/3AsfK7u, and Sinhala writing style classification data set114114114https://bit.ly/3PARuUS) are publicly available. Faumi et al. [136] also worked on Sinhala writing style identification (author identification) using their own data set115115115https://github.com/cipherdragon/SimpleAA which includes Romanized (transliterated) Sinhala text.
The study by Wijayarathna and Jayalal [264] used classical machine learning techniques including Random Forest to classify fake news in Sinhala on Twitter. Chathuranga and Ranathunga [265] used capsule-based methods recommended by Senevirathne et al. [244] to classify Sinhala-English code-mixed data. The work by Weerasiri et al. [266] compared Word2Vec, FastText and Doc2vec [267] against TF-IDF for Sinhala news document classification. The study by Caldera et al. [268] used stacked LSTM to classify Sinhala and Singlish text discussing COVID-19 on Youtube and Twitter.
@@ 3.11 Sentiment Analysis
A simple MLP-based method to classify sentiments in Sinhala text was initially proposed by Medagoda [269] based on their prior work [270]. A word2vec based tool116116116http://bit.ly/2QKI9Np for sentiment analysis of Sinhala news comments is available. A methodology for constructing a sentiment lexicon for Sinhala Language in a semi-automated manner based on a given corpus was proposed by Chathuranga et al. [271]Demotte et al. [272] proposed a sentiment analysis system117117117https://github.com/theisuru/sentiment-tagger based on sentence-state LSTM Networks for Sinhala news comments. In the subsequent work [273], they used word similarity to generate a Sinhala semantic lexicon. They followed this up with a further study [244] which discussed a number of other deep learning models such as RNN and Bi-LSTM in the domain of Sinhala sentiment analysis. Jayasuriya et al. [274] proposed a method to classify Sinhala posts in the domain of sports into positive and negative class sentiments. Ranathunga and Liyanage [275] claimed that using word embedding models as semantic features can compensate for the lack of well developed language-specific linguistic or language resources in the case of analysing sentiment of Sinhala news comments. Jayasuriya et al. [276] conducted a comparative study between word N-grams and character N-grams in the task of semantic classification of Sinhala content in social media. Which they soon followed up with an ensemble approach [277]. The work by Karunarathne [278] used word embedding to analyse the sentiment of manually annotated Sinhala Tweets. Abeyratne and Jayaratne [279] conducted a multi model analysis on classifying Sinhala songs by emotion.
@@
The work by Jayawickrama et al. [280, 281] used the data set released by Wijeratne and de Silva [99] to predict the reactions induced by Sinhala Facebook posts. They then extended the work [282] and compared the results obtained with their data set against that of Senevirathne et al. [244]Aththanayaka and Naleer [283] used Random forest, Support vector machines, and Multinomial Naïve Bayes models to analyse sentiment in Sinhala-English code-mixed text from social media. The ACTSEA dataset118118118https://bit.ly/3oZoziI for Sinhala sentiment analysis was introduced by Jenarthanan et al. [135]. The dataset contains 318,308 Sinhala tweets annotated with emotions. The study by Dhananjaya et al. [284] uses a sentiment lexicon of a high‐resource language to fine-tune Pre‐trained multilingual language models (PMLMs) such as mBERT and XLM‐R on an intermediate task which in turn is then used on the sentiment classification task of Sinhala. A sentiment-annotated dataset of Sinhala-English code mixed YouTube comments has been created by Uthpala and Thirukumaran [285]. However, given that even their paper is behind a paywall, there is no free and public way to access this dataset. The work by Gurgurov et al. [229] compared the performance of LLMs for the Sinhala sentiment analysis task. They report the highest results for the Masked Language Modeling objective for Sinhala in both of their configurations where they used a mixture of ConceptNet (CN) and/or Wikipedia (wiki) language adapters in conjunction with Task Adapters (TA). Liyanaarachchi et al. [286] used character-level embedding to classify sentiments of Sinhala news comments as positive or negative. They claim that the 1-of-m embedding outperforms Log-m embedding for this task. Bandaranayake and Usoof [287] used BERT, DistilBERT, RoBERTa, and XLM-R to analyse the sentiment of Sinhala news comments [244]. They report that XLMR-large model achieves the best result. Fernando and Ranathunga [288] in their study on improving cross-lingual representation of multilingual language models for low-resource languages using linguistic entity masking used Sinhala code mixed sentiment analysis as one of the evaluations.